OpenStudy (anonymous):

Solve for variable: log(x)+log(x+3)=3 log(x^2+3x)=3 10^3=x^2+3x How do you solve for x?

OpenStudy (zzr0ck3r):

10^3 = 1000 = x^2+3x so x^2 + 3x - 1000 = 0

OpenStudy (zzr0ck3r):

(x+(3/2))^2 = 1000 + 9/4 x+(3/2) = +- sqrt(4009/4) x = +- sqrt(4009/4)-(3/2)

OpenStudy (zzr0ck3r):

\[x=-\frac{3}{2}\pm \sqrt{\frac{4009}{4}}\]

OpenStudy (anonymous):

Thank you @zzr0ck3r